Status Report

NASA Space Station On-Orbit Status 11 March 2005

By SpaceRef Editor
March 12, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 11 March 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

After wake-up, morning hygiene and breakfast, the crew worked in the Progress-352 (17P) vehicle, dismantling the Kurs-A rendezvous and approach radar system of its SUDN motion control & navigation system and removing it from the transport drone, a three-hour job. These valuable components will be returned to Earth for reuse. [KURS-A is the active half of the Russian space program’s proven S-band radar system for automated flight, which measures relative motion parameters between Progress (or Soyuz) and the ISS during rendezvous operations, to enable the autopilot’s calculation of corrective impulses. The system’s passive transponder counterpart (KURS-P) is on the Service Module (SM), with one antenna each at the tip of the two solar array wings.]

Later, FE Salizhan Sharipov performed the periodic stand-alone functional closure test of the Vozdukh CO2 removal system’s spare vacuum valve group (BVK) spare. The task was supported by tag-up with ground specialists via S-band. [Closure of the Vozdukh’s vacuum access lines in the event of a malfunction in its BVK vacuum vales or a depressurization in the Vozdukh valve panel (BOA) would be handled by a secondary emergency/backup valve group (AVK). Access to vacuum is required to vent carbon dioxide (CO2) during the regeneration of the absorbent cartridges (PP). During nominal operation, the AVK valves remain open.]

For CDR/SO Chiao, it was the second day of the extensive EVA SPCU (Service & Performance Checkout Unit) repair in the Joint Airlock (A/L), to remove and replace its corroded heat exchanger (HX). [The SPCU HX transfers heat from the EMU/spacesuit water loops to the ITCS LTL (internal thermal control system low temperature loop). Ground testing has shown that the currently installed HX, single-brazed internally with a Nickel (BNi-3) process, could rupture due to BNi-3 corrosion in as early as four years after the Airlock’s launch in July 2001, possibly even earlier. A rupture would be a catastrophic hazard because it could lead to an over-pressurization of the EMU cooling loop, with coolant fluid entering the suit, when the SCU (service &cooling umbilical) is connected to an EMU (the ITCS lines are at ~85 psi while EMU lines are at ~25 psi max.). The new double-brazed HX launched on 17P will have a considerably longer on-orbit life. The HX R&R, spread over several days, started yesterday with video equipment setup to monitor A/L activities, procedures printout, training preps, familiarization with the newly delivered equipment, and tagging up with ground specialists. Yesterday’s scheduled installation of the FSS (Fluid Servicer System) in the A/L for the TCS work was completed today, as were the planned tasks of clearing out A/L stowage, rotating the Cabin Air (CA) rack in the A/L down, out of the way, and disconnecting & isolating TCS loops, with photographs taken. The actual HX R&R will occur on 3/14 (Monday) with leak checks, followed on 3/15 by final ops and rack re-rotation. On 3/16, if everything has gone well, final cleanup, re-stowage, and dismantling of the FSS will conclude the work.]

Sharipov installed the hardware of the European SPQR (Specular Point-like Quick Reference) experiment in the SM, bolting it to window #3 along with its Nikon D1X camera delivered on 17P. The experiment will be conducted from 3/17-3/26 in LVLH XVV attitude in ten sessions, without crew involvement. [SPQR tests a ground-based imaging system, using special optics and image processing, to determine the feasibility of detecting external damage to a spacecraft in orbit from the ground. It uses a Cube Corner Reflector, fixed close to SM Window #3, to reflect a laser beam emitted by a ground station back to the ground. The crew is advised not to look out the ISS windows during the sessions, the times of which will be uplinked. The SPQR Hazard Report indicates that the laser power at the ISS will be well below the threshold for injury.]

Working off the Russian discretionary job jar, Salizhan completed the regular daily inspection of the Lada-5 greenhouse equipment, including filling its water canister as required.

Salizhan completed the routine SOZh/ECLSS servicing/inspection in the SM, including the ASU toilet facilities. The task today also included the installation of an IP-1 air flow sensor at the PrK-SU hatchway from the SM to Progress 17. Later, the FE prepared the daily IMS (Inventory Management System) “delta” file for the automated export/import to the three IMS databases on the ground.

At ~6:00am, Sharipov conducted his regular weekly IMS tagup with specialists at TsUP/Moscow, discussing open issues concerning identification of equipment and storage locations for the IMS databases, via S-band.

The crew performed their regular daily 2.5-hr. physical exercise program on TVIS, RED exerciser, CEVIS cycle and VELO cycle with bungee cord load trainer. [Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the treadmill and one hour on VELO (today: Day 2 of a new set).]

Leroy then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.

At 9:53am, the FE supported a live message/greetings exchange with Valentin Shapiro, CEO of Phonak Co., a manufacturer of high-quality hearing aids, at a commercial PAO event at the European Space Agency’s ESTEC Center in Noordwijk/Holland, requested by Roskosmos/Moscow under contract with ESA. [The event featured the presentation of a pilot hearing aid project called “Savia” that also may become useful on ISS due to its ability to automatically filter out background noise. A backup version of the downlink was taped on 3/7.]

At ~1:30pm EST, the crew held their standard weekly teleconference with the JSC Astronaut Office (Kent Rominger), via S-band S/G, to be followed later, at ~3:45pm, by the regular (nominally weekly) teleconference with the ISS Flight Director at MCC-Houston.

In order to characterize the thermal environment experienced by CMG-1 (control moment gyroscope #1), to be replaced on mission LF-1/STS-114, MCC/H yesterday began an exercise, during which CMG-1 gimbal angles are being commanded to eight different positions, every 12 hours, over the next four days while ISS is still flying in XPOP attitude. Things have been proceeding normally so far. [The same sequence of commanding will also be done in future LVLH XVV and YVV periods prior to the CMG-1 replacement on flight LF-1/STS-114 (switch to LVLH occurs on 3/17). Because the CMG-1 rotor is at standstill, the only heat sources are environmental sources such as solar heating. This thermal mapping may assist in CMG-1 root cause failure analysis after it is brought back on Discovery.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

The Science Officer was thanked by the ADUM (Advanced Diagnostic Ultrasound in Micro-G) team for his feedback at yesterday’s ADUM analysis. [“This communication really helps improve our model for remote guidance, training, and implementation in space flight mission and ground locations. We are looking forward to next month’s scan.”]

Battery reconditioning on set 2B2 continues with no issues so far. Reconditioning of set 2B3 is targeted for April and of 2B1 for early June, after LF-1. [Nickel hydrogen batteries can develop and display “memory loss” resulting in a temporary loss of capacity that should be periodically erased by cycling all material via fully discharging and charging cells (“reconditioning”).]

The Elektron oxygen generator had been functioning nominally in 24 amp mode since 3/9 (Wednesday). Overnight, the primary and secondary pumps failed. The crew was instructed to restart the system in 50 amp mode and then take it to 24 amp mode. Elektron is currently operating nominally. Troubleshooting activities are scheduled for the weekend.

Update on SSC Laptops: Eight A31p NGLs (next generation laptops) are currently in use as SSC machines (1 file server, 6 clients, 1 router). An older IBM 760XD SSC is in use as OCA router, and a 760XD chassis (#6075, former SSC1) is in place as MEC (medical equipment computer).

Next Station Reboost: Currently planned for 3/25 (5:00am EST, 1.65 m/s).

Today’s CEO (crew earth observations) photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in ram), were Tropical Cyclone Willy, Australia (Dynamic Event. A new cyclone has formed to the northwest of Australia and is predicted to reach Category 2 strength at the time of this overpass. Best view of the eye and circulation patterns of the storm was looking back along track), Fires and Smoke, Sumatra (Dynamic Event. A number of large fires are currently burning in central Sumatra and generating extensive smoke plumes. The fires appear to be spaced widely enough that the crew should be able to image the form and direction of individual plumes. This information is useful for determining air circulation patterns associated with the fires), and Mekong River Delta, Vietnam (weather was predicted to be clear for general photography of the Mekong Delta. Upstream changes in land use and coastal development produce rapid and frequent changes to the hydrology and physiographic form of the delta. Wide-field mapping swaths across the delta provide useful context for more detailed photographs).

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Upcoming Key Events:

  • Reboost — 3/25 (5:00am EST; ~1.65 m/s);
  • EVA-13 — 3/25 (could slip to 3/28);
  • Soyuz TMA-6 (10S) launch — 4/15 with Expedition 11 (CDR Sergei Krikalev,
  • FE/SO John Phillips);
  • Soyuz TMA-5 (9S) undocking — 4/25 with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS);
  • LF1 (STS-114) — NET 5/12;
  • Progress M-53 (18P) launch — 6/10;
  • ULF1.1 (STS-121) — NET 7/10;
  • Progress M-54 (19P) launch — 8/24;
  • Soyuz TMA-7 (11S) launch — 9/27.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit (as of this morning, 9:50am EST [= epoch]):

  • Mean altitude — 356.8 km
  • Apogee height — 359.3 km
  • Perigee height — 354.2 km
  • Period — 91.68 min.
  • Inclination (to Equator) — 51.65 deg
  • Eccentricity — 0.0003798
  • Solar Beta Angle — 34.0 deg (magnitude decreasing)
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 160 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 36033

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.